# On the Design and Analysis of LLM-Based Algorithms

### v2

@@ 2.4 Practical considerations
  • Understanding LLM inference service, especially from a system perspective, is crucial for in-depth analysis of cost metrics. LLM inference service can be diverse in practice: for example, LLMs might run on CPUs in a personal laptop or on a distributed GPU cluster, inference might be compute-bound or memory-bound, the complexity of long-sequence processing and generation might be linear or quadratic, parallelism at various levels (e.g. multiple LLMs deployed on multiple machines, or batch inference with one LLM) might be supported, and so on. Fortunately, all these are covered by our proposed framework in a unified manner.
3 Case study: parallel decomposition
This section focuses on parallel decomposition, a basic MapReduce-like [DG08] pattern visualized in Figure 2(a). An algorithm of this pattern first divides the input problem into multiple independent sub-tasks, solves each of them with one LLM node, and finally aggregates the results with an LLM or non-LLM node for the final solution. The intermediate sub-tasks can be solved sequentially or in parallel, which has impacts on certain cost metrics of the overall algorithm. This basic form of task decomposition can be used as a building block for more sophisticated algorithms. Despite its simplicity, interesting analysis and a wide variety of concrete tasks and algorithms can be derived from this pattern, which are presented in the remaining of this section.
3.1 Notations and analysis
Let us start by defining some formal notations that will be useful in our analysis.
  • We denote the size of the input problem instance as $n$. In our study of LLM-based algorithms, the term “size” here typically means the length, e.g. the number of characters or tokens, of the text that represents the input problem instance; it can also be defined in terms of traditional data structures, such as the length of a list.
  • Let $k$ denote the number of parallel sub-tasks after decomposition, and $m_{i}$ denote the size of the $i$-th sub-task, where $i\in[k]$. It is assumed that $m_{i}\leq\overline{m}$ for some maximum value $\overline{m}$ that can fit into the context window of the LLM used in the algorithm. For most of our analysis, we will assume for simplicity that $m_{i}=m$ for all $i\in[k]$, and that $k=O(n/m)$.
  • We denote $p\geq 1$ as the maximum degree of parallelism supported by LLM inference service.
@@ 3.1 Notations and analysis
  • Finally, let $\mathcal{L}_{\textsf{sys}}$ be an upper bound for the length of the system prompt of each LLM call, which includes everything in the prompt except for the size-$n$ input problem instance. The presence of $\mathcal{L}_{\textsf{sys}}$, which can be large in practice, is essentially due to the fact that the LLM is used as a general-purpose sub-routine, hence specifying the concrete task for each LLM call constitutes part of the complexity.
These notations are summarized in Table 1 for convenient reference.
@@
For notational convenience, we will often write $f(n)\lesssim g(n)$ in place of $f(n)=O(g(n))$, which means there exists a universal constant $C>0$ such that $f(n)\leq C\cdot g(n)$ for any positive integer $n$. In addition, $f(n)\asymp g(n)$ means $f(n)\lesssim g(n)$ and $g(n)\lesssim f(n)$ both hold. To further simplify notation, we will often omit the big-O notation in the input arguments of cost functions $\mathcal{C}_{\textsf{pre}}$ and $\mathcal{C}_{\textsf{dec}}$; for example, given a prompt of length $\mathcal{L}_{\textsf{sys}}+O(n)$, we will write the cost of the prefilling phase as $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+n)$ rather than $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+O(n))$.
Analysis of cost metrics.
@@
We assume for simplicity that all parallel LLM nodes share the same LLM model and inference service, and thus the same cost functions $\mathcal{C}_{\textsf{pre}}$ and $\mathcal{C}_{\textsf{dec}}$.
@@
In many cases, the cost of the overall algorithm is a simple sum of the costs of all LLM calls involved. Examples include the financial cost when the LLM APIs are charged by tokens, and the end-to-end latency when all LLM calls are executed sequentially. In such cases, we can write the total cost $\mathcal{C}$ as
@@
[…]
$=k\times\Big{(}\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+m)+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})\Big{)}$
$\lesssim n\times\frac{\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+m)+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})}{m}.$ (4)
Here, the first inequality follows Eq. (2), the second follows $k\lesssim n/m$, and $\mathcal{L}_{\textsf{dec}}=\mathcal{L}_{\textsf{dec}}(m)$ is task-specific, which can be $O(1)$ or $O(m)$ for example.
This basic analysis already provides some hints for tuning the hyperparameter $m$ from the perspective of minimizing costs. Below are a few possible cases, where we assume for simplicity that $\mathcal{L}_{\textsf{dec}}=O(1)$:
[…]
  • It is well known that a Transformer model with full attention suffers from quadratic complexity in long-sequence processing. For this case, a reasonable assumption would be $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+m)+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{sys}}+m,1)\lesssim(\mathcal{L}_{\textsf{sys}}+m)^{2}$, which implies
[…]
    $\lesssim n\times\frac{(\mathcal{L}_{\textsf{sys}}+m)^{2}}{m}=n\times\Big{(}\frac{\mathcal{L}_{\textsf{sys}}^{2}}{m}+m+2\mathcal{L}_{\textsf{sys}}\Big{)}.$
@@
    $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+m)+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{sys}}+m,1)\leq\alpha\times(\mathcal{L}_{\textsf{sys}}+m)^{2}+\beta\times(\mathcal{L}_{\textsf{sys}}+m)+\gamma,$
    which takes into account a more precise characterization of the FLOPs as well as memory IO for loading model weights and KV caches [AQS+23]. In this case, we have
[…]
@@
The above analysis can be extended to the case with parallelism, which is especially relevant to the end-to-end latency of the overall algorithm. Considering parallel LLM calls for homogeneous sub-tasks in an ideal setting where parallelism incurs no additional cost, we have $\mathcal{C}=\mathcal{C}(\text{sub-tasks})+\mathcal{C}(\text{aggregation})$, where
[…]
@@
$\lesssim\frac{n}{p}\times\frac{\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+m)+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})}{m},$
[…]
One implication of the above analysis is that, the optimal value of $m$ that minimizes costs might depend on the choices of cost metrics and assumptions of LLM inference service, among other factors.
Table 1: A list of notations for analysis of parallel decomposition.
Notation Definition
$n$ Size of the input problem instance
$m$ Size of each parallel sub-task, $m\leq\overline{m}$
$k$ Number of parallel sub-tasks, $k\lesssim n/m$
$p$ Maximum degree of parallelism
$\mathcal{L}_{\textsf{sys}}$ Maximum length of the system prompt
Analysis of error metrics.
As explained earlier in Section 2.2, error metrics are mostly task-specific, and thus the analysis of error metrics need to be done in a case-by-case manner. For a given task, error metrics for each homogeneous and parallel sub-task typically increases with the size $m$, although this is not always the case. Error metrics of the overall algorithm after the final aggregation step also depend on the specific task, the method of aggregation, and choices of metrics. We will soon demonstrate the analysis of error metrics in several concrete examples.
3.2 Concrete examples and experiment settings
The rest of this section is dedicated to the analytical and empirical study of a few specific tasks and their corresponding LLM-based algorithms that follow the pattern of parallel decomposition.
Concrete examples.
Tasks that we consider in the rest of this section include counting, sorting, retrieval, retrieval-augmented generation (RAG) and long-text summarization, whose details will soon be explained in their corresponding subsections. For each task, we specify the concrete LLM-based algorithm, analyze its performance in terms of error and cost metrics, and validate our analysis with numerical experiments. Error metrics are task-specific, while cost metrics of interest are common among tasks, including the total prefilling length and decoding length, the total number of LLM calls, and the end-to-end latency with sequential or parallel LLM calls. These concrete examples not only confirm the practical advantages of LLM-based algorithms, but also verify that our analysis can help explain or predict the empirical performance of LLM-based algorithms, reveal the reasons behind some interesting empirical phenomena, and instruct the design of algorithms or choices of hyperparameters, e.g. the sub-task size $m$.

### v3

@@ 2.2 Error and cost analysis
Understanding LLM inference service, especially from a system perspective, is crucial for in-depth analysis of cost metrics. LLM inference service can be diverse in practice: LLMs might run on CPUs in a personal laptop or on a distributed GPU cluster, inference might be compute-bound or memory-bound, the complexity of long-sequence processing and generation might be linear or quadratic, parallelism at various levels (e.g., multiple LLMs deployed on multiple machines, or batch inference with one LLM) might be supported, and so on. These are covered by our proposed framework in a unified manner.
3 Case studies and insights: parallel decomposition
This section is dedicated to parallel decomposition, a basic MapReduce-like pattern (Dean & Ghemawat, 2008) visualized in Figure 3(a). For concreteness, we consider three canonical tasks, which are abstract and synthetic versions of practical LLM applications. (1) Counting: given a string of length $n$ consisting of letters and digits, count the number of digits in it. (2) Sorting: given a list of $n$ numbers, sort it in ascending order. (3) Retrieval: answer a given question that requires retrieving some key information from a long piece of text. Table 1 summarizes the LLM-based algorithms for solving these tasks by parallel decomposition, and Table 2 presents a worked example for the counting task. Supplementary details for these tasks, as well as additional examples of parallel decomposition, can be found in Appendix A.
A problem of particular interest in practice is: what is the proper granularity of task decomposition, or the sub-task size $m$, for solving a task of size $n$ by parallel decomposition? More specifically, we strive to answer the following questions:
Q1. How does the overall cost of the LLM-based algorithm change with the sub-task size $m$? Q2. Does a smaller $m$, i.e., finer-grained task decomposition, always imply a smaller final error?
Throughout this section, we answer these questions by leveraging the analytical framework proposed in Section 2. The results are summarized in Insights 1 and 2, which will also be validated empirically.
Why synthetic tasks?
For the purpose of building a solid foundation for rigorously understanding the accuracy and efficiency of LLM-based algorithms, we have chosen to stick with synthetic settings for our empirical studies in this work. This brings various benefits, such as reducing the risk of test data contamination, facilitating unambiguous and precise measurement of accuracy, allowing full transparency and control over task configurations, and making the current work as self-contained as possible. We start with the simplest settings and algorithms (for didactic purpose) in this section, and will gradually expand our proposed framework to more complex and practical scenarios resembling those in prior empirical works, e.g., retrieval-augmented generation in Appendix A.5, iterative retrieval and reasoning in Section 4.2, or active information seeking and aggregation with recursive task decomposition in Section 4.3.
Formal notations.
@@ 3 Case studies and insights: parallel decomposition
We let $n$ denote the size of the input problem instance; it can be the length (e.g., the number of characters or tokens) of the text that represents the input problem instance, or defined in terms of traditional data structures (e.g., the length of a list). Let $k$ denote the number of parallel sub-tasks after decomposition, and $m_{i}$ denote the size of the $i$-th sub-task, where $i\in[k]$. It is assumed that $m_{i}\leq\overline{m}$ for some maximum value $\overline{m}$ that can fit into the context window of the LLM used by the algorithm. For most of our analysis, we will assume for simplicity that $m_{i}=m$ for all $i\in[k]$, and that $k=\Theta(n/m)$. We denote $p\geq 1$ as the maximum degree of parallelism supported by LLM inference service. Finally, let $\mathcal{L}_{\textsf{sys}}$ be an upper bound for the length of the system prompt of each LLM call, which includes everything in the prompt except for the input problem instance itself. The presence of $\mathcal{L}_{\textsf{sys}}$, which can be large in practice, is essentially due to the fact that the LLM is used as a general-purpose sub-routine, hence specifying the concrete task for each LLM call constitutes part of the complexity.
@@
For notational convenience, we will often write $f(n)\lesssim g(n)$ in place of $f(n)=O(g(n))$, which means there exists a universal constant $C>0$ such that $f(n)\leq C\cdot g(n)$ for any positive integer $n$. In addition, $f(n)\asymp g(n)$ means $f(n)\lesssim g(n)$ and $g(n)\lesssim f(n)$ both hold. To further simplify notation, we will often omit the big-O notation in the input arguments of cost functions $\mathcal{C}_{\textsf{pre}}$$\mathcal{C}_{\textsf{dec}}$$\mathcal{C}_{\textsf{LLM}}$; for example, given a prompt of length $\mathcal{L}_{\textsf{sys}}+O(n)$, we will write the cost of the prefilling phase as $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+n)$ rather than $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+O(n))$.
Table 1: Three concrete examples of parallel decomposition.
Task Counting Sorting Retrieval
Step 1 Divide the input string into $k$ disjoint sub-strings of lengths $m_{1},m_{2},\dots,m_{k}$. Divide the input list into $k$ disjoint sub-lists of lengths $m_{1},m_{2},\dots,m_{k}$. Divide the input text into $k$ overlapping chunks of lengths $m_{1},m_{2},\dots,m_{k}$.
Step 2 For each $i\in[k]$, use one LLM call to count the number of digits in the $i$-th sub-string, which returns an answer $y_{i}$. For each $i\in[k]$, use one LLM call to sort the $i$-th sub-list, which returns a solution $\bm{y}_{i}$. For each $i\in[k]$, use one LLM call to try to answer the question based on that chunk.
Step 3 The final solution of the algorithm is $y=\sum_{i\in[k]}y_{i}$. Merge the sub-lists $\bm{y}_{1},\dots,\bm{y}_{k}$ into a single list $\bm{y}$ using a symbolic algorithm (Listing 1). Generate the final answer by majority voting of the valid answers.
We adopt majority voting because, as the input text is chunked with overlapping, the correct answer might appear multiple times in Step 2. Moreover, there might exist false positives in Step 2, in which case majority voting brings extra robustness.
In Step 2, an LLM call might return “I don’t know” if it decides that the corresponding chunk does not contain sufficient information for answering the question. Such an answer is considered invalid, and excluded from majority voting in Step 3.
Table 2: A worked example for the counting task, showing the prompt and response of a specific LLM node.
1# Prompt
2User: Count the number of digits in the following string:
3dhd73695051jcbcg9jbeh903bi64ic736icfag96
4Answer in the following format: There are {answer} digits in the given string’.
5Do NOT output anything else.
6
7# Response
8Assistant: There are 19 digits in the given string
3.1 Answering Q1: cost analysis
@@
We assume for simplicity that all parallel LLM nodes share the same LLM backend and inference service, and thus the same cost functions $\mathcal{C}_{\textsf{pre}}$, $\mathcal{C}_{\textsf{dec}}$ and $\mathcal{C}_{\textsf{LLM}}$.
The direct sum of costs.
@@
In many cases, the cost of the overall algorithm is the direct sum of the costs of all LLM calls involved, e.g., the monetary cost of API calls for proprietary LLMs, or the end-to-end latency when all LLM calls are executed sequentially. In such cases, we can write the total cost $\mathcal{C}$ as
@@
$\mathcal{C}(\text{sub-tasks})$ $=k\times\mathcal{C}(\text{one sub-task})\leq k\times\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})$
$\lesssim n\times\frac{\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})}{m}.$ (4)
Here, the first inequality follows Eq. (2), the second inequality follows $k\lesssim n/m$, and $\mathcal{L}_{\textsf{dec}}=\mathcal{L}_{\textsf{dec}}(m)$ is task-specific, e.g., $O(1)$ for the counting and retrieval tasks, or $O(m)$ for the sorting task.
This basic analysis already provides some hints for tuning the hyperparameter $m$:
  • If $\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})=\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+m)+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})$ grows with $m$ at a linear or sub-linear rate, then the right-hand side of Eq. (4) is monotonely decreasing in $m$, which means $\mathcal{C}(\text{sub-tasks})$ is minimized at $m=\min\{n,\overline{m}\}$.
  • A Transformer model with full attention suffers from quadratic complexity in long-sequence processing (Yuan et al., 2024; Zhou et al., 2024). With the simplified assumption that $\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})\lesssim(\mathcal{L}_{\textsf{sys}}+m)^{2}$, we have
    $\mathcal{C}(\text{sub-tasks})$ $\lesssim n\times\frac{\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})}{m}\lesssim n\times\frac{(\mathcal{L}_{\textsf{sys}}+m)^{2}}{m}=n\times\Big(\frac{\mathcal{L}_{\textsf{sys}}^{2}}{m}+m+2\mathcal{L}_{\textsf{sys}}\Big).$
@@
    $\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})\leq\alpha\times(\mathcal{L}_{\textsf{sys}}+m)^{2}+\beta\times(\mathcal{L}_{\textsf{sys}}+m)+\gamma,$
    which takes into account a more precise characterization of LLM inference service, including the FLOPs as well as memory IO for loading model weights and KV caches. In this case, we have
    $\mathcal{C}(\text{sub-tasks})$ $\lesssim n\times\frac{\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})}{m}$
@@
The above analysis can be extended to the case with parallelism, which is especially relevant to the end-to-end latency of the overall algorithm. Considering parallel LLM calls for homogeneous sub-tasks in an ideal setting where parallelism incurs no additional cost, we have
@@
$\mathcal{C}(\text{sub-tasks})$ $\leq\Big\lceil\frac{k}{p}\Big\rceil\times\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})$ (5)
according to Eq. (3). For large $m$ such that $k\leq p$, we have $\lceil k/p\rceil=1$, and thus $\mathcal{C}(\text{sub-tasks})\leq\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})$ is monotonely increasing in $m$. On the other hand, for sufficiently small $m$ and hence large $k$, we have $\lceil k/p\rceil\approx k/p$, and thus the upper bound for $\mathcal{C}(\text{sub-tasks})$ can be approximated by
$\mathcal{C}(\text{sub-tasks})$ $\leq\frac{k}{p}\times\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})\lesssim\frac{n}{p}\times\frac{\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})}{m},$
which might be monotonely decreasing in $m$ if $\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})$ grows with $m$ at a linear or sub-linear rate. In this case, the overall cost $\mathcal{C}(\text{sub-tasks})$ is minimized by $k\asymp p$ and hence $m\asymp n/p$.
Summary.
With the above analysis, our answer to Q1 is summarized as follows:
Insight 1.
The relationship between the overall cost $\mathcal{C}(\text{sub-tasks})$ and the sub-task size $m$ depends on various factors, such as the choices of cost metrics and assumptions about LLM inference service. For example: (1) the direct sum of costs is minimized at a smaller value of $m$ if each sub-task cost $\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}}(m))$ grows with $m$ at a faster rate; (2) if $\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}}(m))$ grows with $m$ (sub-)linearly, then the end-to-end latency with parallelism degree $p$ is minimized at $m\asymp n/p$.
3.2 Answering Q2: error analysis
While error analysis is task-specific and has to be conducted case by case, we will see that general principles can be derived for the pattern of parallel decomposition, to be summarized in Insight 2.
3.2.1 Counting
Assume for notational convenience that $m_{i}=m$ for all $i\in[k]$, and $k=n/m$ is an integer. We denote the ground-truth count for the complete string as $y^{\star}$, and the ground-truth count for the $i$-th sub-string as $y_{i}^{\star}$.
If $\mathcal{E}$ represent the absolute counting error, then the overall error is bounded by the sum of sub-task errors:
If $\mathcal{E}$ represent the normalized counting error instead, then
$\mathcal{E}(y)\coloneqq\frac{|y-y^{\star}|}{n}\leq\frac{1}{n}\sum_{i\in[k]}|y_{i}-y_{i}^{\star}|=\frac{1}{k}\sum_{i\in[k]}\frac{|y_{i}-y_{i}^{\star}|}{m}=\frac{1}{k}\sum_{i\in[k]}\mathcal{E}(y_{i}),$
i.e., the overall error is bounded by the average of sub-task errors; since a smaller value of $m$ makes each sub-task easier, it is reasonable to expect that the overall error $\mathcal{E}(y)$ will also become smaller as $m$ decreases.
3.2.2 Sorting
Compared to counting, there are more diverse phenomena in the sorting task in terms of error metrics. In particular, there exist various failure modes in sorting with an LLM-based algorithm: (1) the output list might not be monotone; (2) the length of the output list might be larger or smaller than that of the input list; (3) the output list might contain numbers that do not match exactly those of the input list.
To accommodate these failure modes, we may define the following error metrics for sorting a list, where $\bm{y}$ denotes the solution returned by the algorithm and $\bm{y}^{\star}$ denotes the ground-truth solution:
Note that the same error metrics can be similarly defined for each parallel sub-task in Step 2 of the LLM-based algorithm, and it is reasonable to expect that they become smaller as the sub-task size $m$ decreases. On the other hand, analyzing the errors of the overall algorithm after the final merging step can be more complicated. As an example, focusing on the third failure mode and the $\ell_{\infty}$ error metric, we have the following theoretical guarantee (whose proof is deferred to Appendix A.3.4):
Proposition 1.
Consequently, if each sub-task error is monotone in $m$, then this error bound is also monotone in $m$.
3.2.3 Retrieval
For the retrieval (or needle-in-a-haystack (Kamradt, 2023)) task, we start by identifying two failure modes of each LLM call for retrieving the targeted information from a chunk of size $m$ in Step 2 of the algorithm:
  1. 1.
    The LLM might mistakenly return “I don’t know” or an incorrect passcode when the chunk contains the specific needle that is necessary and sufficient for answering the question. In general, this failure mode will occur more frequently for larger values of $m$. Our early experiments with various LLMs confirmed that this failure mode begin to manifest when $m$ exceeds a certain LLM-specific threshold.
  2. 2.
    The LLM might mistakenly return a wrong answer when the needle is actually absent from the chunk (i.e., false positives). We observed empirically that some LLMs are more susceptible to this failure mode even when the value of $m$ is small, while others are less so.
Based on the above observations, we can hypothetically categorize LLMs into two types: Type-A LLMs are only prone to the first failure mode, while Type-B LLMs are prone to both.
It turns out that error analysis for the overall LLM-based algorithm depends critically on which type of LLM is used. If a Type-A LLM is used, a smaller value of $m$ means the first failure mode is less likely to occur in Step 2 of the algorithm, which implies higher accuracy for the final solution of the overall algorithm. On the other hand, analysis is more complicated if a Type-B LLM is used, since both failure modes can possibly occur in Step 2 of the algorithm: a larger value of $m$ means the first failure mode is more likely to occur, while a smaller value of $m$ implies a larger number of chunks $k\asymp n/m$, which can potentially increase the chance of error in the final step of the algorithm (i.e., majority voting), due to the frequent occurrence of the second failure mode in Step 2 of the algorithm. Consequently, the minimum error of the overall algorithm might be achieved by some intermediate value of $m$ that achieves a balance between the two failure modes. If $n$ is too large, then there might not exist a good value of $m$ that can achieve a low error, as either failure mode must occur with high probability.